OpenStudy (anonymous):

Write a ratio for each word phrase in reduced form. 48 inches to 5 feet.

OpenStudy (jdoe0001):

how many inches in 1 foot anyway?

OpenStudy (anonymous):

i think 12

OpenStudy (jdoe0001):

so in 5 feet, there'll be 12 + 12 + 12 + 12 + 12 = 60 or 5 * 12 = 60 inches then

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ 5\cdot \cfrac{12in}{ft}\implies 60in}} \\ \quad \\ \quad \\ \textit{48 inches to 5 feet}\implies \textit{48 inches to 60 inches} \\ \quad \\ \implies \cfrac{48}{60}\Leftarrow\textit{divide top and bottom by 12}\)

OpenStudy (jdoe0001):

well... \(\bf {\color{blue}{ 5\cancel{ft}\cdot \cfrac{12in}{\cancel{ft}}\implies 60in}} \\ \quad \\ \quad \\ \textit{48 inches to 5 feet}\implies \textit{48 inches to 60 inches} \\ \quad \\ \implies \cfrac{48}{60}\Leftarrow\textit{divide top and bottom by 12}\)

OpenStudy (anonymous):

4/5

OpenStudy (jdoe0001):

yeap, then the reduced form will 4 to 5, or 4:5